# Thomason condition for regular algebraic stacks

### v2

@@ Appendix A Descending Thomason condition
Next, we check that $\mathbf{R}f_{\ast}D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})\subseteq D_{\operatorname{qc},Z}(\mathcal{X})$. However, being that $D_{\operatorname{qc}}(\mathcal{Z})$ is compactly generated by the collection $B\otimes^{\mathbf{L}}P$ with $B\in\mathcal{B}$, every object of $D_{\operatorname{qc}}(\mathcal{Z})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $B\otimes^{\mathbf{L}}P$ (see [Sta25, Tag 09SN]). So, this can be checked by showing $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)\in D_{\operatorname{qc},Z}(\mathcal{X})$. However, this can be seen by reducing to schemes (see e.g. [GW23, Remark 23.46(2)]).
@@
Now, we finish the proof. As we have shown the adjunction restricts, we can make use of the fact that $f^{!}$ is conservative by [HR17a, Theorem 4.14(4)]. Consequently, Lemma˜4.2 implies the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. This tells us every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ (see [Sta25, Tag 09SN]). So, if the support of every $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ is properly contained in $Z$, then so must the support of every object in $D_{\operatorname{qc},Z}(\mathcal{X})$, which is absurd. Therefore, there must exist some $B\in\mathcal{B}$ such that $\operatorname{supp}(\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P))=Z$. ∎

### v3

@@ 4. Proofs
Next, we check that $\mathbf{R}f_{\ast}D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})\subseteq D_{\operatorname{qc},Z}(\mathcal{X})$. Recall that $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ is compactly generated by the collection $B\otimes^{\mathbf{L}}P$ with $B\in\mathcal{B}$. Hence, every object of $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $B\otimes^{\mathbf{L}}P$ (see [STA26, Tag 09SN]). So, this can be checked by showing $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)\in D_{\operatorname{qc},Z}(\mathcal{X})$. However, this can be seen by reducing to schemes (see e.g. [GW23, Remark 23.46(2)]).
@@
Now, we finish the proof. As we have shown the adjunction restricts, we can make use of the fact that $f^{\times}$ is conservative by [HR17a, Theorem 4.14(4)]. Consequently, Corollary˜3.6 implies the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. This tells us every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of the collection $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ (see [STA26, Tag 09SN]). Hence, if the support of every $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ is properly contained in $Z$, then so must the support of every object in $D_{\operatorname{qc},Z}(\mathcal{X})$, which is absurd. Therefore, there must exist some $B\in\mathcal{B}$ such that $\operatorname{supp}(\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P))=Z$. ∎
Secondly, we need a version of [NEE23, Observation 5.6] which is applicable to quasi-affine monomorphisms, which appears in the proof of Theorem˜1.1.
Lemma 4.4.
Let $t\colon\mathcal{Y}\to\mathcal{X}$ be a quasi-affine morphism of quasi-compact quasi-separated algebraic stacks. Suppose $t$ is injective on the underlying topological spaces. If $\mathcal{X}$ satisfies the $\beta$-Thomason condition, then so does $\mathcal{Y}$.
Proof.
As $t$ is quasi-affine, it must be concentrated. So, [HR17a, Lemma 8.2] guarantees that $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by a collection of at most $\leq\beta$ objects. Next, let $Z$ be a closed subset of $|\mathcal{Y}|$. Set $Z^{\prime}$ to be the closure of $t(Z)$ in $|\mathcal{X}|$. Then $t^{-1}(Z^{\prime})=Z$ because $t$ is injective on the underlying topological spaces. Moreover, the hypothesis on $\mathcal{X}$ allows us to find a $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z^{\prime}$. However, [HR17a, Lemma 4.8(2)] tells us $t^{-1}(\operatorname{supp}(P))=\operatorname{supp}(\mathbf{L}t^{\ast}P)$, which completes the proof because $\mathbf{L}t^{\ast}P\in\operatorname{Perf}(\mathcal{Y})$. ∎
Now, we prove the main result.
Proof of Theorem˜1.1.
The proof proceeds by an inductive argument. However, to set the stage, we start with a construction needed for such an argument. By [HR18, Theorem 4.1], there exist morphisms of algebraic stacks $V\xrightarrow{p}\mathcal{Y}\xrightarrow{f}\mathcal{X}$ satisfying
  • $V$ is an affine scheme
  • $p$ is finite, flat, surjective, and of finite presentation
  • $f$ is a Nisnevich covering (see [HR18, Definition 3.1]) of finite presentation with separated diagonal (in fact, $f$ is étale and surjective).
Moreover, from [HR18, Proposition 3.1], there exists a monomorphic splitting sequence for $f$ in the sense of loc. cit. Specfically, we have a sequence of open immersions
such that the base change of $f$ to $|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|$ (when given the reduced closed substack structure) admits a monomorphic section for each $c=1,\ldots,n$.
Next, before going to the induction argument, we make an observation for components of the construction above. Fix $c=1,\ldots,n$. Consider the fibered square
$\mathcal{Y}_{c}$$(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}$$\mathcal{Y}$$\mathcal{X}$$\scriptstyle{f_{c}}$$\scriptstyle{h_{c}}$$\scriptstyle{j^{\prime}_{c}\circ t_{c}}$$\scriptstyle{f}$
obtained by base change of $f$ along $j^{\prime}_{c}\circ t_{c}$ where $j^{\prime}_{c}$ is the associated open immersion and $t_{c}$ the associated closed immersion. Choose a monomorphic section $g_{c}\colon(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}\to\mathcal{Y}_{c}$. This gives us a commutative diagram
[…]
Base change tells us that $f_{c}$ is étale and of finite presentation. Hence, $g_{c}$ must be étale (see e.g. [STA26, Tag 0CIR]). Furthermore, since $g_{c}$ is monomorphic, it is separated, of finite type, and representable by algebraic spaces (see e.g. [STA26, Tag 06MY & Tag 050J]). Yet, each algebraic stack in our proof is Noetherian, so $g_{c}$ must be of finite presentation (see e.g. [STA26, Tag 0DQJ]). It follows that $g_{c}$ must be quasi-affine (use e.g. [OS03, Proposition 3.1]). There is a fibered square,
[…]
Again, from base change, we know that $h_{c}$, and hence $h^{\prime}_{c}$, is an immersion. Since any such morphism is quasi-affine, it follows that $V_{c}$ is a quasi-affine scheme. Hence, by Proposition˜4.3, we know that $\mathcal{Y}_{c}$ must be $1$-Thomason because $p_{c}$ is a finite, flat, and surjective morphism of finite presentation from a quasi-affine scheme. Thus, from Lemma˜4.4, we know that each $(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}$ must be $1$-Thomason because $g_{c}$ is a quasi-affine monomorphism morphism (see e.g. [STA26, Tag 0500]). Consequently, as each $\mathcal{X}_{c}$ is regular, Lemma˜4.1 tells us that $D_{\operatorname{qc},Z_{c}}(\mathcal{X_{c}})$ is singly compactly generated where $Z_{c}:=t_{c}(|(|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|)_{red}|)$. So, we may find a $P_{c}\in\operatorname{Perf}(\mathcal{X}_{c})$ which compactly generates $D_{\operatorname{qc},Z_{c}}(\mathcal{X_{c}})$.
Now, we prove the desired claim by induction on $n$. Specifically, we are inducting the length of the monomorphic splitting sequence of $f$. There is nothing to check if $n=0$. Moreover, the observation above addresses the base case $n=1$. So, we may assume that $n\geq 2$. By [HLL+25, Proposition B.1], there is a Verdier localization sequence
$D^{b}_{\operatorname{coh},Z_{c}}(\mathcal{X}_{c})\xrightarrow{(i_{c})_{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{X}_{c})\xrightarrow{\mathbf{L}j_{c-1}^{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{X}_{c-1})$
where $(i_{c})_{\ast}$ is the natural inclusion. However, [DLM+25a, Theorem 3.7] tells us that $\operatorname{Perf}=D^{b}_{\operatorname{coh}}$ in each case. This allows us to find a $G_{c}\in\operatorname{Perf}(\mathcal{X}_{c})$ such that $\mathbf{L}j_{c-1}^{\ast}G_{c}$ generates $D_{\operatorname{qc}}(\mathcal{X}_{c-1})$ and belongs to $\operatorname{Perf}(\mathcal{X}_{c-1})$. Consider the recollement obtained in Proposition˜3.1,
$D_{\operatorname{qc}}(\mathcal{X}_{c-1})$$D_{\operatorname{qc}}(\mathcal{X}_{c})$$D_{\operatorname{qc},Z_{c}}(\mathcal{X}_{c})$$\scriptstyle{\mathbf{R}(j_{c-1})_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}_{c-1}}$$\scriptstyle{j^{\times}_{c}}$$\scriptstyle{i^{!}_{c}}$$\scriptstyle{(i_{c})_{\ast}}$$\scriptstyle{i_{c}^{\ast}}$
where $(i_{c})_{\ast}$ is the natural inclusion and $j^{\times}_{c}$ the right adjoint of $\mathbf{R}j_{\ast}$. So, in particular, $(i_{c})_{\ast}P_{c}=P_{c}$. Then, from Proposition˜3.7, we know that $G_{c}\oplus P_{c}$ generates $D_{\operatorname{qc}}(\mathcal{X}_{c})$. Consequently, this establishes the induction step. Since $f$ always has monomorphic splitting sequence of finite length, we complete the proof. ∎
Remark 4.5.
In the proof of Theorem˜1.1, if $G_{c}$ can be chosen to be compact, then it follows that $D_{\operatorname{qc}}(\mathcal{X}_{c})$ is singly compactly generated. However, it is not clear to us whether this is always possible.
Lemma 4.6.
Let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack. Suppose $\mathcal{S}\subseteq\operatorname{Perf}(\mathcal{X})$ generates $D_{\operatorname{qc}}(\mathcal{X})$. Consider a quasi-compact open immersion $j\colon\mathcal{U}\to\mathcal{X}$ with complement $Z:=|\mathcal{X}|\setminus|\mathcal{U}|$. If there is a $Q\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(Q)=Z$, then $D_{\operatorname{qc},Z}(\mathcal{X})$ is generated by the collection $S\otimes^{\mathbf{L}}Q$ where $S\in\mathcal{S}$.
Proof.
This mimics the proof of [HR17a, Lemma 4.10(2)], but without compactness. Suppose $E\in D_{\operatorname{qc},Z}(\mathcal{X})$ satisfies $\operatorname{Hom}(S\otimes^{\mathbf{L}}Q[n],E)\cong 0$ for all $S\in\mathcal{S}$ and $n\in\mathbb{Z}$. By adjunction, it follows that $\operatorname{Hom}(S[n],\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(Q,E))\cong 0$. However, $\mathcal{S}$ generates $D_{\operatorname{qc}}(\mathcal{X})$, and so $\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(Q,E)\cong 0$. Since $\operatorname{supp}(Q)=Z$, it follows from [HR17a, Lemma 4.9] that $E\cong 0$, which completes the proof. ∎
Corollary 4.7.
Let $\mathcal{X}$ be a quasi-compact quasi-separated regular algebraic stack with quasi-finite diagonal. Then $D_{\operatorname{qc},Z}(\mathcal{X})$ is generated by a single perfect complex for all $Z=|\mathcal{X}|\setminus|\mathcal{U}|$ with $\mathcal{U}\to\mathcal{X}$ a quasi-compact open immersion.
Proof.
Since $\mathcal{X}$ is regular and Noetherian, we know that $\operatorname{Perf}(\mathcal{X})=D^{b}_{\operatorname{coh}}(\mathcal{X})$ (see e.g. [DLM+25a, Theorem 3.7]). Denote by $i\colon\mathcal{Z}\to\mathcal{X}$ for the closed immersion from the reduced induced closed substack structure on $Z$. Hence, $\mathbf{R}i_{\ast}\mathcal{O}_{\mathcal{Z}}\in\operatorname{Perf}(\mathcal{X})$ with $\operatorname{supp}(\mathbf{R}i_{\ast}\mathcal{O}_{\mathcal{Z}})=Z$. Consequently, the desired claim follows from Theorem˜1.1 and Lemma˜4.6. ∎
Proof of Corollary˜1.2.
If $\mathcal{X}$ is concentrated, then $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$. So, the desired claim follows from Corollary˜4.7 with $Z=|\mathcal{X}|$. In fact, this shows that $\mathcal{X}$ satisfies the $1$-Thomason condition, see Remark˜3.2. ∎